OpenStudy (anonymous):

Solve the equations x-y-5=0, x^2+2xy+y^2 . . . ANSWER={(4,-1),(1,-4)}

OpenStudy (anonymous):

can u complete the question

OpenStudy (anonymous):

x^2+2xy+y^2=?????

OpenStudy (anonymous):

X^2+2XY+Y^2=9

OpenStudy (anonymous):

First we have to eliminate one variable: x-y=5 -->(A) (x+y)^2= (x-y)^2+4xy=25+4xy=9 substitute x-y=5 in above, we get xy=(9-25)/4=-4 => y=-4/x substitute this in (A), we get x+4/x = 5 x^2+4-5x=0 Now solve this to find x, and from (A) u can find y

OpenStudy (anonymous):

X^2+2XY+Y^2=9 or (x+y)^2=9 thus taking sqaure toots u will have x+y=+3(eq2) or x+y=-3(eq3) let x-y=5 be eq1 now u have to solve two pairs simultaneously equations viz 1 ....is x-y-5=0, and x+y=3 and 2.... x-y-5=0, and x+y=-3 now for 1 add eq1 and eq2 we have 2x=8 or x=4 hence putting x=4 in x+y=3 we have y=-1 thus (4,-1) is one of the answer again for 2 add eq1 and eq3 we have 2x=2 or x=1 hence putting x=1 in x+y=-3 we have y=-4 thus (1,-4) is one of the answer thus the answers are (4,-1) and (1,-4)